OpenStudy (ebydaqueen19):

help me plzzzz!!! this is my equation: CaO + H2SO4 > CaSO4 + H20 Calculate the mass of CaSO4 that can be produced when 4.070 g of CaO react with excess sulfuric acid.

OpenStudy (ebydaqueen19):

is this a percent yield problem?

OpenStudy (anonymous):

\[Ca O+ H _{2}S O _{4}=Ca SO4+H _{2}\] 40+16 +......=40+32+4*16+.... 56+...=136+... 56 gms .of CaO produce CaSO4=136 gms 1 gm. of CaO produce Ca S O4=136/56 gms. 4.070 gms. of CaO produce Ca SO4\[=\frac{ 136 }{ 56 }\times 4.070=?\]

OpenStudy (ebydaqueen19):

9.884285714 thts what i got

OpenStudy (ebydaqueen19):

i only need 2 significant figures right....since the 56 o0nly has 2

OpenStudy (anonymous):

i think it should be 9.884 up to three decimal places if your calculations are correct.

OpenStudy (ebydaqueen19):

oh ok thnks

OpenStudy (anonymous):

yw